11x + 7 -4=10x - 3
Answer by ReadingBoosters(3246)

Starting with
11x + 7 - 4 = 10x - 3
Combine like terms
11x - 10x + 3 - 3 = 10x - 10x - 3 - 3
x = -6
Proof
11(-6) + 7 - 4 = -66 + 7 - 4 = -63
10(-6) - 3 = -60 - 3 = -63